OpenStudy (anonymous):

A rectangular barge 6 m long and 3 m wide floats in fresh water. How much will the barge sink when loaded with 600 kg of sand? Show all calculations leading to an answer.

OpenStudy (jjuden):

600kg sand displaces 600kg water. Water's density is ≈ 1,000kg/m³ The displaced water occupies about 600/1000=.6m³ The barge's area is 15m². V = L×W×H = .6m³ H (depth of sinking) = .6m³/15m² = .04m = 40mm

OpenStudy (anonymous):

@jjuden could you explain how you knew water displaces equally with sand